It works like this: each number on a die can generate 6 combinations. Two can generate $6^2$ or $36$ combinations. Why? Write out all scenarios
- (1,1), (1,2), (1,3), (1,4), (1,5), and (1,6)
- (2,1), (2,2), (2,3), (2,4), (2,5), and (2,6)
- (3,1), (3,2), (3,3), (3,4), (3,5), and (3,6)
- (4,1), (4,2), (4,3), (4,4), (4,5), and (4,6)
- (5,1), (5,2), (5,3), (5,4), (5,5), and (5,6)
- (6,1), (6,2), (6,3), (6,4), (6,5), and (6,6)
See? 36 possible combinations. Now, see how (1,1) is just 1 of the 36. This means there's a 1 in 36 chance of getting (1,1) or $1/36=0.027$ or $2.7\%$.
Rememeber also our theme here: "believable" probability always requires a large number of trials. Think of this when you choose your
N.
Any thoughts on any other two combinations the dice could show? (Like 5 and 7?)